OpenStudy (anonymous):

Find the value of c such that the system of equations cx-16y=-5 9x-12y=-11 does not have a solution.

OpenStudy (mertsj):

Remember that parallel lines do not intersect and so a system of parallel lines has no solution. Remember also that parallel lines have the same slope. What is the slope of the second line?

OpenStudy (anonymous):

9/12

OpenStudy (anonymous):

3/4

OpenStudy (mertsj):

So make c a number that will cause the slope of the first line to also be 3/4

OpenStudy (anonymous):

umm no idea

OpenStudy (anonymous):

i dont really need to know this, we just have to do it for the class, so can you just tell me the answer?

OpenStudy (anonymous):

please (:

OpenStudy (mertsj):

How did you get the slope of the second line?

OpenStudy (anonymous):

y=mx+b

OpenStudy (mertsj):

Do that to the first line.

OpenStudy (anonymous):

1/16?

OpenStudy (mertsj):

cx-16y=-5 -16y=-cx-5 \[y=\frac{-c}{-16}-5\] \[y=\frac{c}{16}-5\]

OpenStudy (mertsj):

Make that +5/16

OpenStudy (mertsj):

\[y=\frac{c}{16}x+\frac{5}{16}\]

OpenStudy (mertsj):

So we see that the slope is c/16 which much be 3/4

OpenStudy (mertsj):

And c/16=3/4 So what is c?

OpenStudy (anonymous):

12?

OpenStudy (mertsj):

Yes. If c =12, the lines are parallel and the system has no solution.

OpenStudy (anonymous):

THANKS! i have two more, could you help me with those too? please??(:

OpenStudy (anonymous):

-9x-15y=-11 bx+20y=10

OpenStudy (mertsj):

Same thing.

OpenStudy (mertsj):

Try it again and I will check your work.

OpenStudy (anonymous):

y=-3/5

OpenStudy (mertsj):

Yes that is the slope of the first line.

OpenStudy (anonymous):

-b/20 is the second one?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

does b=10?

OpenStudy (mertsj):

\[\frac{-b}{20}=\frac{-3}{5}\]

OpenStudy (mertsj):

\[-5b=-60\]

OpenStudy (mertsj):

b=12

OpenStudy (anonymous):

okay thank you so much! and the last one is 12x+dy=-3 -8x+10y=-3

OpenStudy (mertsj):

Same thing.

OpenStudy (anonymous):

so the slope of the first one is 4/5

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

the second one y=-12/d - 3?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

d=18?

OpenStudy (mertsj):

Actually -3/d but it doesn't affect this problem

OpenStudy (anonymous):

wait nevermind lol

OpenStudy (mertsj):

no

OpenStudy (anonymous):

-12/d=4/5?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so confused

OpenStudy (mertsj):

Cross multiply.

OpenStudy (anonymous):

15?

OpenStudy (mertsj):

\[\frac{-12}{d}=\frac{4}{5}\] 4d=-60 d=-15

OpenStudy (anonymous):

YAY THANK YOU

OpenStudy (mertsj):

yw